# Quasi-stationary evolution of cubic-quintic NLSE drop-like solitons in DNA-protein systems

### v1

@@ Abstract
Nonlinear molecular excitations in DNA have traditionally been modelled using the nonlinear Schrödinger equation (NLSE). An alternative approach is based on the generalized plane-base rotator model and the $SU(2)/U(1)$ generalized spin coherent states, which leads to a cubic-quintic NLSE. The higher-order nonlinearities are particularly useful for modelling complex interactions, such as those in DNA-protein systems, where multiple competing forces play a significant role. Additionally, the surrounding viscous medium introduces dissipative forces that affect the propagation of molecular excitations, leading to energy dissipation and damping effects. These damping effects are modelled using the quasi-stationary method, which describes the system’s near-equilibrium behaviour. In this work, we explore the evolution of nonlinear molecular excitations in DNA-protein systems, accounting for damping effects, and discuss potential applications to the transcription process.
@@ 1 Introduction
Understanding the physical characteristics of genomic DNA, such as base pair deformations, impurities, and domain walls, is essential for comprehending its structure and recognizing specific sequences [1]. These features not only influence DNA stability and functional properties but also play a critical role in protein binding and gene regulation [3, 4, 5, 6, 7]. Consequently, new methods are needed to analyze and accurately represent the sequence-dependent bending and twisting of neighboring base pairs [8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. However, the sequence-specific recognition of DNA by proteins and other ligands demands a more sophisticated description of DNA’s flexibility than can be provided by traditional uniform elastic models. DNA undergoes sequence-dependent conformational changes, such as kinking and intercalation, in response to protein binding, making it necessary to account for the dynamic nature of these interactions [19, 20]. Moreover, both DNA and its binding proteins experience conformational changes as they form functional complexes and bind with other molecules. This is true for processes like chromosome condensation and segregation, which depend on DNA-binding proteins that facilitate chromosome organization through DNA bridging and bending [21]. These conformational changes have been explained theoretically in terms of long-lived and large amplitude localized nonlinear excitations. More broadly, it is well-established that such excitations play a key role in a variety of biological processes involving DNA, including transcription, gene expression, and the recognition of promoter sites, among others [22]. Experimentally, ensemble approaches including surface plasmon resonance, white light interferometry and electrophoretic mobility shift assay are capable of measuring protein-DNA binding but are not sensitive to how these proteins change DNA conformation [18]. Exploring these conformational changes requires a more controlled environment, which has recently been provided by DNA-designed crystals [23, 24, 25, 26, 27, 28]. This approach enables precise manipulation and analysis of DNA structural features, interactions, and behaviours, thereby enhancing our understanding of its dynamics and stability. As this field continues to grow, it is expected to yield new insights into DNA-protein dynamics in the near future [29].
@@
As can be seen, the study of DNA conformational changes resulting from interactions with other molecules remains a long-standing challenge. However, several efforts have been made to investigate these phenomena through the development of mathematical models. In this context, the mathematical model of double-stranded DNA proposed by Takeno and Homma [30, 31], commonly known as the plane base-rotator model, has gained widespread acceptance. Using a generalized version of this model: M. Daniel and his collaborators devoted a series of studies to the internal dynamics of inhomogeneous double-stranded DNA molecular system, which dynamics have been governed by the sine-Gordon equation and the nonlinear schrödinger equation [32, 33, 34, 35]; Agüero et al, using the generalized coherent states approach in Perelomov sense studied the DNA-protein interaction obtaining several soliton-like solutions contained on the classical motion equations [36, 37]; Saha and Kofané employed the plane base-rotator model to study numerically the effects of inhomogeneities on the double stranded DNA interacting with RNAp enzyme [38, 39]; Saccomandi studied the inhomogeneous stacking on the double strand using the Peyrard-Bishop model and Takeno-Homma model [40]; and so forth.
@@
In most of the studies mentioned above, the internal dynamics of base pairs, influenced by protein interactions, were described using nonlinear Schrödinger equations with higher-order nonlinearities. This kind of equations have widely extended in many branches of mathematical physics in order to explain nonlinear phenomena like water waves or the propagation of light pulses in fibers for example [41, 42, 43, 44, 45, 46]. A very particular case of interest is the nonlinear Schrödinger equation with cubic-quintic nonlinearities commonly known as the cubic-quintic NLSE, which appears in many branches of physics. For instance in: the nuclear hydrodynamics with Skyrme forces [47]; the Bose-Einstein condensates [48, 49]; plasma physics [50]; generalized elastic solids [51]; and the principal applications are in nonlinear optics for the description of the optical pulse propagations in dielectric media of non-Kerr type [52].
Another important contribution to a more accurate description of the DNA-protein system and its associated nonlinear excitations is the inclusion of a viscous medium as a damping term. The addition of this term, mathematically leads to non-integrability of the resulting equation. Thus, exact $N$- soliton solutions cannot be obtained, and a perturbative analysis becomes necessary. This issue is particularly evident in the nonlinear Schrödinger equation framework [32]. In the case of the NLSE with higher-order nonlinearities, the non-integrability is less obvious, since Lax pairs cannot be directly constructed, and thus no $N$-soliton solutions can be derived. Such equations arise when the generalized coherent states approach is employed, leading to a cubic-quintic NLSE. Consequently, studying the adiabatic evolution of solitons is crucial for understanding the dynamics of DNA base pairs.
@@
The aim of the present work is to investigate the evolution of nonlinear excitations in the DNA-protein molecular system embedded in a viscous medium. This analysis was performed using a multi-scale approach, which allows for the identification of localized perturbations around the soliton. Through this method, we observed how the soliton parameters are modified by the perturbative term. A similar study was conducted in the context of optical fibers [53], but the approach can also be extended to biological systems, such as lipid membranes in the framework of a Boussinesq-like equation [54]. Motivated by these previous works, we present the present analysis, which is organized as follows: In Section 2, we provide a brief overview of the plane base-rotator model, where the dynamics of the DNA base pairs with a binding protein are described by the equations of motion derived from the generalized Hamiltonian. After some simplifications, these equations are reduced to the nonlinear Schrödinger equation with cubic-quintic nonlinearities, facilitated by the use of $SU(2)/U(1)$ generalized coherent states. Section 3 focuses on the interaction between the DNA and the surrounding medium. As is well known, the viscous medium introduces dissipative forces that affect the propagation of molecular excitations, leading to energy dissipation and damping effects. To model the steady-state or near-equilibrium behavior of the system, we adopt the quasi-stationary approximation, in which the system evolves slowly relative to the dissipation timescale. This approximation is applied to the cubic-quintic NLSE. In Section 4, we apply these analytical results to the transcription process. Finally, Section 5 concludes the paper with a summary and discussion of our findings.
@@ 2.1 Plane-base rotator model of a DNA-protein system
Consider the well-known B-form of DNA, where the sugar-phosphate backbone is represented by two ribbons, $S$ and $S^{\prime}$, that spiral around each other along the $z$-axis, as illustrated in Fig. 1a [30, 31]. The coordinates $P_{n}$ and $P_{n}^{\prime}$ represent the points where the $n$-th and $n^{\prime}$-th bases bases attached to their respective strands. These points are given by $(r\cos(n\varphi_{0}),r\sin(n\varphi_{0}),z_{n})$ and $(r\cos(n\varphi_{0}+\pi),r\sin(n\varphi_{0}+\pi),z_{n}^{\prime})$, respectively, where $\varphi_{0}=2\pi/p$ with $p=10$ being the number of bases per turn in $S$ and $S^{\prime}$. Here, $r$ is the radius of the circle, as shown in Fig. 1b. Additionally, $(\theta_{n},\varphi_{n})$ and $(\theta_{n}^{\prime},\varphi_{n}^{\prime})$ represent the angles of rotation of the $n$-th base pair around the points $P_{n}$ and $P_{n}^{\prime}$, respectively.
@@
Below, we outline the general procedure for deriving the generalized Hamiltonian as an extension of the plane-base rotator model (for further details, see [33, 36]).
@@
    The hydrogen bonds energy, or inter-strand energy, can be derived from the heuristic argument that it depends on the distance between the tip of the $n$-th base and its complementary $n^{\prime}$-th base base in the opposing strand, denoted as $\overline{Q_{n}Q_{n}^{\prime}}$. For simplicity, we assume that the positions of the bases , $z_{n}=z_{n}^{\prime}$, are identical, as shown in Fig. 1c.
@@
    The quasi-spin operators $\vec{S}_{n}=(S_{n}^{x},S_{n}^{y},S_{n}^{z})$, expressed in terms of the rotational angles $\theta$ and $\varphi$, with components
    $S_{n}^{x}=\sin\theta_{n}\cos\varphi_{n},\quad S_{n}^{x}=\sin\theta_{n}\sin\varphi_{n}\quad\text{and}\quad S_{n}^{x}=\cos\theta_{n}$
@@
    The kinetic and potential energies of DNA and protein are added, where $X_{n}$ stands for the $n$-th base displacement along the direction of the hydrogen bond, and $Y_{n}$ denotes the displacement for $n$-th peptide in the protein chain, with $p_{n}$ and $q_{n}$ as corresponding momenta.
@@
Considering all the listed above requirements, it is obtained the following Hamiltonian
@@
here, $J$ and $\mu$ are constants representing the stacking interaction energy between the $n$-th base and its nearest neighbours in the two strands, and the hydrogen bond energy between the complementary bases, respectively. Additionally, $k_{1}$ and $k_{2}$ are elastic constants that characterize the behaviour of the hydrogen atom attached to the base and the amplitude of oscillations in the protein. The masses $m_{1}$ and $m_{2}$ correspond to the hydrogen atom attached to the base and the peptide in the protein, respectively. Finally, $\alpha_{1}$ and $\alpha_{2}$ are coupling coefficients that describe the interaction between thermal phonons and the DNA and protein, respectively
@@ 2.2 Weakly saturable approximation and cubic-quintic NLSE
$S^{+}=\langle\psi\rvert\hat{S}^{+}\lvert\psi\rangle=\dfrac{\bar{\psi}}{1+|\psi|^{2}};\quad S^{-}=\langle\psi\rvert\hat{S}^{-}\lvert\psi\rangle=\dfrac{{\psi}}{1+|\psi|^{2}};\quad S^{z}=\langle\psi\rvert\hat{S}^{z}\lvert\psi\rangle=-\dfrac{1-|\psi|^{2}}{2(1+|\psi|^{2})}$ (2)
@@
similarly, for $X_{n+1}$, $Y_{n+1}$, $\xi_{n+1}$ and $|\xi_{n\pm 1}|$. This is possible because the length of excitations in DNA is much greater than the inter-site distance $a$ between neighbouring nucleotides. Therefore, by averaging the Hamiltonian (1) using the expressions in (2) and incorporating the above considerations, we obtain:
@@
Thus, the classical motion equations for $\psi$, $\xi$, where by symmetry of the double-stranded DNA $\xi=-\psi$, as well as for the displacements $X$ and $Y$, can be derived directly from (5) in a similar way. After reducing the system of coupled equations (for further details and references, see [36]), we obtain
@@
here, we consider the parametric change $z=s\sqrt{\frac{Ja}{2}}$, along with the condition that the parameter $\mu$ satisfy the strong inequality $\frac{\mu}{a}\gg 1$. Additionally, we define the following parameters
$\beta=\frac{\alpha_{1}}{2(m_{1}v_{1}^{2}-2k_{1})}\quad\text{and}\quad\gamma=\frac{\alpha_{2}}{2(m_{2}v_{2}^{2}-2k_{2})}.$ (7)
Expression (6) represents a generalized nonlinear Schrödinger equation with saturable nonlinearities, which fully describes the dynamics of the DNA-protein system. The analytical solutions of generalized nonlinear Schrödinger equations in saturable media present an intriguing challenge, and while some attempts have been made to solve the simplest versions of this equation directly [58, 59], a general solution remains elusive. In future work, we plan to explore the analytical solutions of Eq. (6). For the present study, however, we employ a weakly saturable approximation, where $G(I)=F(I)(1+I)^{-1}\approx F(I)(1-I)$ with $I=|\psi|^{2}$. After performing the reparametrization $\mu/a\to\mu$, we obtain the cubic-quintic NLSE, which is given by
$i\psi_{t}+\psi_{ss}+k_{1}\psi+k_{3}|\psi|^{2}\psi+k_{5}|\psi|^{4}\psi=0.$ (8)
@@
$k_{1}=-2\mu+\beta+\gamma,\quad k_{3}=4\mu-10\beta-6\gamma,\quad k_{5}=4\mu-34\beta-18\gamma$ (9)
In order to analyze the effects of the viscous medium on the system, we consider a simpler form of Eq. (8)
$i\varphi_{t}+\frac{1}{2}\varphi_{ss}+{\varphi}{|\varphi|^{2}}-\alpha\varphi{|\varphi|^{4}}=0.$ (10)
which was obtained by the following scale transformation
$\psi=e^{ik_{1}t}\phi(s,t),\ \phi=\varphi(\beta s,\beta^{2}{t})$
where $\beta=(k_{3}/2)^{1/2}$ and $\alpha=-2k_{5}/k_{3}$. Henceforth, we consider a reparametrizated form of Eq. (10), that is $s\to x$.

### v2

@@ Abstract
Nonlinear molecular excitations in DNA have traditionally been studied within the framework of the nonlinear Schrödinger equation (NLSE). An alternative approach is based on the plane-base rotator model and $SU(2)/U(1)$ generalized spin coherent states, which lead to a cubic-quintic NLSE. Higher-order nonlinearities are particularly useful for modelling complex interactions, such as those in DNA-protein systems, where multiple competing forces play a significant role. Additionally, the surrounding viscous medium introduces dissipative forces that influence the propagation of molecular excitations, resulting in energy dissipation and damping effects. These damping effects are modelled using the quasi-stationary method, which describes the system’s near-equilibrium behaviour. In this work, we analyse the evolution of nonlinear molecular excitations in DNA-protein systems, accounting for damping effects, and discuss potential applications to the transcription process.
@@ 1 Introduction
The local physical properties of genomic DNA, including base pair deformations, impurities, and domain walls, are crucial in regulating stability, gene expression, sequence recognition, and protein binding [1, 2, 4, 5, 6, 7, 8]. For example, DNA undergoes sequence-dependent conformational changes, such as kinking and intercalation, in response to protein binding [9, 10]. This is particularly important in processes like chromosome condensation and segregation, which rely on DNA-binding proteins that facilitate chromosome organization through DNA bridging and bending [11]. Thus, the sequence-specific recognition of DNA by proteins and other ligands requires a more nuanced understanding of DNA’s flexibility—an aspect that traditional uniform elastic models fail to fully capture. Therefore, accurately describing, analyzing, and representing these conformational changes demand the development of new ones [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. These conformational changes can be depicted as long-lived and large amplitude localized nonlinear excitations. More broadly, it is well-established that such excitations play a key role in a variety of biological processes involving DNA, including transcription, gene expression, and the recognition of promoter sites, among others [22]. Experimentally, ensemble techniques such as surface plasmon resonance, white light interferometry, and electrophoretic mobility shift assays are commonly used to measure protein-DNA binding. However, these methods do not provide insight into how proteins induce conformational changes in the DNA [23]. To investigate these structural changes, a more controlled environment is required, and DNA-designed crystals have emerged as a promising tool for this purpose [24, 25, 26, 27, 28, 29]. This innovative approach enables a detailed analysis of DNA’s structural and physical properties, along with its interactions and behaviour, thereby advancing our understanding of the macromolecule. As this field continues to grow, it is expected to yield new insights into DNA-protein dynamics soon [30].
@@
Despite significant efforts, studying DNA conformational changes resulting from interactions with other molecules remains a long-standing challenge. However, this can be addressed through the implementation of mathematical models. In this context, the double-stranded model of DNA, introduced by Takeno and Homma [31, 32] and commonly referred to as the plane base-rotator model, has become widely accepted. Using a generalized version of this model: M. Daniel and his collaborators devoted a series of studies to the internal dynamics of inhomogeneous double-stranded DNA molecular system, which dynamics have been governed by the sine-Gordon equation and the NLSE [33, 34, 35, 36, 37]; Agüero et al, using the generalized coherent states approach, in Perelomov sense, studied the DNA-protein interaction obtaining several classical and non-classical soliton solutions that describe the open states of base pairs [38, 39, 40]; Saha and Kofané employed the plane base-rotator model to study numerically the effects of inhomogeneities on the double stranded DNA interacting with RNAp enzyme [41, 42]; Saccomandi studied the inhomogeneous stacking on the double strand using the Peyrard-Bishop model and Takeno-Homma model [43].
@@
In most of the studies mentioned above, the internal dynamics of base pairs, influenced by protein interactions, were described using NLSE with different nonlinearities. This kind of equation have widely extended in many branches of mathematical physics in order to explain nonlinear phenomena like water waves or the propagation of light pulses in fibres for example [44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. A very particular case of interest is the NLSE with cubic-quintic nonlinearities commonly known as the cubic-quintic NLSE, which appears in many branches of physics ranging from Bose-Einstein condensates to the description of optical pulse propagations in dielectric media of non-Kerr type, including plasma physics, nuclear hydrodynamics, generalized elastic solids, among others [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
@@
Another important contribution to a more accurate description of the DNA-protein system and its associated nonlinear excitations is the inclusion of a viscous medium as a damping term. The addition of this term mathematically leads to non-integrability of the resulting equation. Thus, exact $N$- soliton solutions cannot be obtained, and a perturbative analysis becomes necessary [33]. In the case of the NLSE with higher-order nonlinearities, the non-integrability is less obvious, since Lax pairs cannot be directly constructed, and thus no $N$-soliton solutions can be derived. Such equations arise when the generalized coherent states approach is employed, leading to a cubic-quintic NLSE. The aim of the present work is to investigate the evolution of nonlinear excitations in the DNA-protein molecular system embedded in a viscous medium. This analysis is performed using a multi-scale perturbation approach, which allows for the identification of localized perturbations around the soliton. Through this method, we observed how the soliton parameters are modified by the perturbative term. A similar study was conducted in the context of optical fibres [79], but the approach can also be extended to biological systems, such as lipid membranes in the framework of a Boussinesq-like equation [80]. Motivated by these previous works, we present this analysis, which is organized as follows: In Section 2, we provide a brief overview of the plane base-rotator model, where the dynamics of the DNA base pairs with a binding protein are described by the equations of motion derived from the generalized Hamiltonian. After some simplifications, these equations are reduced to the NLSE with cubic-quintic nonlinearities, facilitated by the use of $SU(2)/U(1)$ generalized coherent states. Section 3 focuses on the interaction between the DNA and the surrounding medium. As is well known, the viscous medium introduces dissipative forces that affect the propagation of molecular excitations, leading to energy dissipation and damping effects. To model the steady-state or near-equilibrium behavior of the system, we adopt the quasi-stationary approximation, in which the system evolves slowly relative to the dissipation timescale. In Section 4, we apply these analytical results to the transcription process. Finally, Section 5 concludes the paper with a summary and discussion of our findings.
@@ 2.1 Plane-base rotator model of a DNA-protein system
Let us consider the B-form of DNA, where the sugar-phosphate backbone is represented by two ribbons, $S$ and $S^{\prime}$, that spiral around each other along the $z$-axis, as illustrated in Fig. 1a [31, 32]. The coordinates $P_{n}$ and $P_{n}^{\prime}$ represent the points where the $n$-th and $n^{\prime}$-th bases are attached to their respective strands. These points are given by $(r\cos(n\varphi_{0}),r\sin(n\varphi_{0}),z_{n})$ and $(r\cos(n\varphi_{0}+\pi),r\sin(n\varphi_{0}+\pi),z_{n}^{\prime})$, where $\varphi_{0}=2\pi/p$. Here, $r$ is the radius of the circle and $p=10$ is the number of bases per turn in $S$ and $S^{\prime}$, respectively, as shown in Fig. 1b. Additionally, $(\theta_{n},\varphi_{n})$ and $(\theta_{n}^{\prime},\varphi_{n}^{\prime})$ represent the angles of rotation of the $n$-th and $n^{\prime}$-th base pairs around the points $P_{n}$ and $P_{n}^{\prime}$, correspondingly.
@@
Below, we outline the general procedure for deriving the generalized Hamiltonian as an extension of the original plane-base rotator model (for further details, see [35, 38]).
@@
    The hydrogen bond (or inter-strand) energy can be derived from the heuristic argument that it depends on the distance between the tip of the $n$-th base and its complementary $n^{\prime}$-th base in the opposing strand, denoted as $\overline{Q_{n}Q_{n}^{\prime}}$. For simplicity, we assume that the positions of the bases, $z_{n}=z_{n}^{\prime}$, are identical, as shown in Fig. 1c.
@@
    The quasi-spin operators $\vec{S}_{n}=(S_{n}^{x},S_{n}^{y},S_{n}^{z})$, expressed in terms of the rotational angles $\theta_{n}$ and $\varphi_{n}$, with components
    $S_{n}^{x}=\sin\theta_{n}\cos\varphi_{n},\quad S_{n}^{y}=\sin\theta_{n}\sin\varphi_{n}\quad\text{and}\quad S_{n}^{z}=\cos\theta_{n}$
@@
    The kinetic and potential energies of the DNA, along with the protein energy, are summed up, where $X_{n}$ stands for the $n$-th base displacement along the direction of the hydrogen bond, and $Y_{n}$ denotes the displacement for $n$-th peptide in the protein chain. The corresponding momenta are $p_{n}$ and $q_{n}$, respectively.
@@
Taking into account all the aforementioned requirements, the following Hamiltonian is obtained:
@@
here, $J$ and $\mu$ are constants that represent the stacking interaction energy between the $n$-th base and its nearest neighbours in the two strands, and the hydrogen bond energy between the complementary bases, correspondingly. Additionally, $k_{1}$ and $k_{2}$ are elastic constants that characterize the behaviour of the hydrogen atom attached to the base and the amplitude of oscillations in the protein. The masses $m_{1}$ and $m_{2}$ correspond to the hydrogen atom attached to the base and the peptide in the protein, respectively. Finally, $\alpha_{1}$ and $\alpha_{2}$ are coupling coefficients that describe the interaction between thermal phonons, DNA and the protein.
@@ 2.2 Weakly saturable approximation and cubic-quintic NLSE
$S^{+}=\langle\psi\rvert\hat{S}^{+}\lvert\psi\rangle=\dfrac{\bar{\psi}}{1+|\psi|^{2}};\quad S^{-}=\langle\psi\rvert\hat{S}^{-}\lvert\psi\rangle=\dfrac{{\psi}}{1+|\psi|^{2}};\quad S^{z}=\langle\psi\rvert\hat{S}^{z}\lvert\psi\rangle=-\dfrac{1-|\psi|^{2}}{2(1+|\psi|^{2})},$ (2)
@@
similarly, for $X_{n+1}$, $Y_{n+1}$, $\xi_{n+1}$ and $|\xi_{n\pm 1}|$. This is possible once that the length of excitations in DNA is much greater than the inter-site distance $a$ between neighbouring nucleotides. Therefore, by averaging the Hamiltonian (1) using the expressions in (2) and incorporating the above considerations, we obtain:
@@
Thus, the classical motion equations for the hydrogen atom displacement and the peptide displacement turns out to be
$X_{z}=\dfrac{\alpha_{1}}{2(m_{1}v_{1}-2k_{1})}\dfrac{1+|\psi|^{4}-6|\psi|^{2}}{(1+|\psi|^{2})^{2}}$ (6)
and
$Y_{z}=\dfrac{\alpha_{2}}{2(m_{2}v_{2}-2k_{2})}\left(\dfrac{1-|\psi|^{2}}{1+|\psi|^{2}}\right)^{2}$ (7)
with the governing equation for $\psi$ being
@@
here, we consider the parametric change $z=s\sqrt{\frac{Ja}{2}}$, along with the condition $\frac{\mu}{a}\gg 1$.
Additionally, we define the following parameters
$\beta=\frac{\alpha_{1}^{2}}{2(m_{1}v_{1}^{2}-2k_{1})}\quad\text{and}\quad\gamma=\frac{\alpha_{2}^{2}}{2(m_{2}v_{2}^{2}-2k_{2})},$ (9)
with $v_{1}$ and $v_{2}$ as velocities of the hydrogen atom and the peptide, respectively. In order to obtain (8) we need to consider the symmetry of the base pairs in the complementary strands, i.e., $\psi=-\xi$ as well as Eqs. (6) and Eq. (7) (additional details can be found in [38]). Equation (8) represents a generalized NLSE with saturable nonlinearities, which fully describes the dynamics of the DNA-protein system. The analytical solutions of the generalized NLSE in saturable media represent an intriguing challenge, and while some attempts have been made to solve the simplest versions of this equation directly [84, 85], a general solution remains elusive. For the current study, however, we employ a weakly saturable approximation, where $G(I)=F(I)(1+I)^{-1}\approx F(I)(1-I)$ with $I=|\psi|^{2}$. After performing the reparametrization $\mu/a\to\mu$, we obtain the cubic-quintic NLSE, which is given by
$i\psi_{t}+\psi_{ss}+k_{1}\psi+k_{3}|\psi|^{2}\psi+k_{5}|\psi|^{4}\psi=0,$ (10)
@@
$k_{1}=-2\mu+\beta+\gamma,\quad k_{3}=4\mu-10\beta-6\gamma,\quad k_{5}=4\mu-34\beta-18\gamma.$ (11)
It is evident that the use of generalized coherent states associated with the $SU(2)/U(1)$ group provides a more general framework than that derived from the standard coherent states. This assertion can be readily demonstrated by showing that, with a careful choice of parameters, we can reduce our governing equation to previously studied cases, as we explore in the following limiting cases.
  1. 1.
    By selecting the parameters $\mu$, $\beta$ and $\gamma$ such that $k_{5}=0$, we find that Eq. (10) simplifies to the NLSE, which has been already obtained in the previously mentioned studies of open states of the base pairs [33, 34, 35, 36, 37]. As it is well known the trivial boundary condition $x\to\pm\infty$ and $\psi\to 0$, with $k_{3}>0$ and normalized constants, lead to the a collective wave of bell type which form an open state with for the angle deviation, as indicated by Eq. (3). This wave is expected to propagate along the DNA molecule due to interactions with thermal phonons surrounding the molecule. The collective wave is constructed from two types of bell-shaped solutions for the $\psi$ and $\xi$ fields, respectively.
  2. 2.
    Another limiting case of Eq. (10) arises when $k_{3}=0$. By selecting appropriate values of $\mu$, $\beta$, and $\gamma$ and $k_{5}>0$, we obtain the attractive Schrödinger equation with quintic nonlinearity or quintic Schrödinger equation. In this case, the evolution equation splits into two sectors. The first sector corresponds to collapse, where the amplitude of the solutions grows infinitely in a finite time. The second sector represents dispersion, where the amplitude decays and the localized region expands. The first sector holds no physical significance; therefore, effects that counteract this collapse must be introduced, such as inelastic collisions of DNA base pairs, which would lead to energy loss in the form of heat emission.
As shown earlier, our governing equation can indeed be directly reduced to the well-known cases. However, it also provides additional physical insights beyond what is already known. In addition to the classical soliton structures that form open states, as derived from the previously discussed cases, we also encounter non-classical soliton structures, such as compactons and anti-compactons. The compacton solution appears as a result of the interaction of two drop solitons for $\psi$ and $\xi$, once we replace the solutions in the expression for the hydrogen atom $X$ given by Eq. (6) with $\psi$ being the solution of Eq. (8) with the boundary conditions $\psi\to 0$ as $x\to\pm\infty$. It’s worth to point out that when the two solutions travel along the DNA chain the compacton forms open states and the anti-compacton breaks this deformation. Thus, the anticompactons play a major role in the DNA repair mechanism.
Both classical and non-classical structures are highly sensitive to medium effects, which can substantially alter their profiles. As such, studying these structures is crucial for understanding the dynamics of open states in DNA chains, especially in relation to protein interactions.